OpenStudy (anonymous):

when a 140g piece of metal at 70 degree Celsius is placed in 220g of water at 0 degree Celsius, the metal is cooled, and the water is warmed and both come to a final temperature of 10 degree Celsius. What is the specific heat of the metal?

OpenStudy (owlfred):

Hoot! You just asked your first question! Hang tight while I find people to answer it for you. You can thank people who give you good answers by clicking the 'Good Answer' button on the right!

OpenStudy (anonymous):

wow, this one is thermodynamics right?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

its actually physics but it's considered to be in the math category so i need a lot of help with this.

OpenStudy (anonymous):

The heat associated with a certain mass changing temperature is

OpenStudy (anonymous):

m*c*(dT)

OpenStudy (anonymous):

the energy change in the water has to equal the energy change in the metal

OpenStudy (anonymous):

Qm=140*c*(70-10) Qw=220*4.186*(10-0)

OpenStudy (anonymous):

you set the two equal to each other and solve for c. That equation will give units of Joule/(gram Celsius)

OpenStudy (anonymous):

so would i be solving the equation that you suggested i'm confused

OpenStudy (anonymous):

Have you learned the heat/energy required to change the temperature of a substance?

OpenStudy (anonymous):

no i don't quite understand how to solve problems like these.

OpenStudy (anonymous):

ok we have the equation Q=mc(dT) where Q is heat, m is mass of whatever is changing temperature, dT is the change in Temperature.

OpenStudy (anonymous):

the magnitude of the heat has to be the same for both the metal and the water

OpenStudy (anonymous):

so you plug the known values into the heat equation for water and metal and set them equal to each other to find the specific heat you are looking for.

OpenStudy (anonymous):

The magnitude of the heat is the same in both due to conservation of energy

OpenStudy (anonymous):

so basically what would be the numbers in the equation to plug into the problem if is the same.

OpenStudy (anonymous):

Your final equation would look like: 140*c*(70-10)=220*4.186*(10-0) 140*60*c=220*10*4.186

OpenStudy (anonymous):

and you want to find c in J/gC

OpenStudy (anonymous):

so how would i find c

OpenStudy (anonymous):

140*60*c=220*10*4.186 8400*c=9209.2 c=9209.2/(8400) c=1.1 J/gC

OpenStudy (anonymous):

1.0963 would be c

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

so what units should i use for this problem?

OpenStudy (anonymous):

that c is J/(gC) or Joules per gram degree Celsius

OpenStudy (anonymous):

the energy in Joules it takes to change the temperature of 1 gram of metal 1 degree celsius

OpenStudy (anonymous):

so would this be the same as 1.0963 cal/g degree Celsius ?

OpenStudy (anonymous):

no 1 cal= 4.186 Joules

OpenStudy (anonymous):

if you want calories divide your answer by 4.186

OpenStudy (anonymous):

i don't think my answer is right

OpenStudy (anonymous):

i have 0.2619

OpenStudy (anonymous):

yep I think that's good

OpenStudy (anonymous):

why dont you think it is right

OpenStudy (anonymous):

i don't know because i think that my units wont match up with the right answer that i just put up there

OpenStudy (anonymous):

im not sure what you mean

OpenStudy (anonymous):

like should it say 0.2619 cal/g degrees Celsius, does that sound right?

OpenStudy (anonymous):

oh yeah those are the units for specific heat

OpenStudy (anonymous):

thank you so much I got it right thanks

OpenStudy (anonymous):

you're welcome :)